OpenStudy (anonymous):

There are 12 students in a social studies class. Three students will be selected to present their term projects today. In how many specific ways can three students be selected?

OpenStudy (anonymous):

\[_{12}C_3\] is what you are being asked for

OpenStudy (luigi0210):

\[\frac{ 12! }{ (12-3)! }\]

OpenStudy (anonymous):

12/9?

OpenStudy (anonymous):

That's not combination. By specific way do we mean that ORDER is important? That decides combination or permutation.

OpenStudy (luigi0210):

It does say specific

OpenStudy (anonymous):

A. 1,320 B. 220 C. 504 D. 36 Are my choices...

OpenStudy (anonymous):

You are on track with my thoughts Luigi. Even if we pick random kids, they are 1st, 2nd, and 3rd in presentation.

OpenStudy (anonymous):

Combinations without Repetition \[C(n.r) = _{n}C _{r} = \frac{ n! }{ r!(n-r)! }\] ---------------------------------------------- Permutations without Repetition \[P(n.r) = _{n}P _{r} = \frac{ n! }{ (n-r)! }\]

OpenStudy (anonymous):

Very nice presentation some_someone!

OpenStudy (anonymous):

Oh! I see.

OpenStudy (anonymous):

12C3=220?

OpenStudy (luigi0210):

So it is combinations..?

OpenStudy (anonymous):

yep its a combination

OpenStudy (anonymous):

That's where the 12 and 3 go, now it makes sense.

OpenStudy (anonymous):

I wish I could give medals to all of you.

OpenStudy (anonymous):

well then give it to the one who you think helped you the most :)